The lenght breadth height of a cuboid block are in the ratio 5:4:3 its surface area is1216cmsq what is the volume of the solid
1
Expert's answer
2020-03-31T09:22:04-0400

Let the length, breadth and height of the cuboid be 5x, 4x and 3x, respectively.


"A=2(ab+bc+ac)=2(20+12+15)x^2=1216"

"94x^2=1216"

"x=4\sqrt{\frac{38}{47}}"

"V=abc=(5)(4)(3)\left(4\sqrt{\frac{38}{47}}\right)^3"

"V=2792\ cm^3"